-1/3 divide 2/5 = 25/9 divide x
Answer by rfer(16322)

(-1/3)/(2/5)=(25/9)/(x)
-1/3x=2/5*25/9
-1/3x=50/45
-1/3x=10/9
x=(10/9)/(-1/3)
x=10/9*-3/1
x=-30/9
x=-10/3